What is the inflation rate between 2015 & 2016?

Prices Quantities
2015 2016 2015 2016
Soda $2 $2 36 40
Burgers $8 $12 8 10
French Fries $2 $1 16 20
Answer #1

Cost of market basket in 2015 = ($2 * 36) + ($8 * 8) + ($2 * 16)

= 72 + 64 + 32 = $168

Cost of market basket in 2016 =($2 * 40)+($12 * 10)+($1 * 20)

= 80 + 120 + 20 = $220

CPI = (current year cost of market basket / base year cost of market basket) * 100

CPI in 2015 = (168 / 168) * 100 = 100

CPI in 2016 = (220 / 168) * 100 = 130.95

Inflation rate between 2015 & 2016

= [(130.95 - 100) / 100] * 100

= 30.95%